OpenStudy (smurfy14):

P+Dgy+(1/2)D(v^2)=C Solve for D.

OpenStudy (lalaly):

\[D(gy+\frac{v^2}{2})=C-P\]\[D=\frac{C-P}{gy+\frac{v^2}{2}}\]